OpenStudy (anonymous):

Find the derivative of x/ radical (x^4 + 4) Please help!!!

OpenStudy (anonymous):

You'll want to use the quotient rule with f(x) = x and g(x) = √(x^4 + 4). To find g'(x) just use the chain rule with f(x) = √g(x) and g(x) = x^4 + 4.

OpenStudy (anonymous):

so is this the first step. [-1/2(x^4 + 4)^-3/2 * 4x^3] [x] - [1] [radical(x^4 + 4)] / [radical (x^4 + 4)^2]

OpenStudy (anonymous):

@Fruitbasket i'm having trouble simplifying.

OpenStudy (anonymous):

I think there's something wrong with that expression. \[y = \frac{ x }{ \sqrt{x^4 + 4}}\] With the quotient rule, you take the x on top as f(x) (which you did) but you take √(x^4 + 4) as g(x), not 1/√(x^4 + 4). Otherwise, your method is correct.

OpenStudy (anonymous):

i multiplied it by one. i didn't divide it

OpenStudy (anonymous):

I mean on the bit inside the first bracket, you got -1/2(x^4 + 4)^-3/2 by differentiating (x^4 + 4)^-1/2. It should be 1/2(x^4 + 4)^-1/2 from differentiating (x^4 + 4)^1/2.